OpenStudy (anonymous):

the length of a puppy on a scale drawing is 6 cm. if the scale is 1/9 cm, what is the actual length of the puppy? a.14 cm b.54 cm c.48 cm d.15 cm

OpenStudy (anonymous):

6cm=1/9 of actually size. So 6*9=?cm = answer.

OpenStudy (_.mufasa._):

B. 54 cm

OpenStudy (anonymous):

ok stay i need to ask you another

OpenStudy (anonymous):

the scale on a drawing of a swimming pool is 1 cm=4m. what is the actual length of the pool if it is 10 cm long on the drawing? a.40 m b.14 m c.40 cm d.15 cm

OpenStudy (anonymous):

4*10=?

OpenStudy (_.mufasa._):

C. 40 cm

OpenStudy (anonymous):

ok i got 2 more

OpenStudy (anonymous):

the distance between morristown and bellington is 4 inches on a map. If the scale on the map is 1 inch=25 miles, what is the actual distance between the two towns? a.29 miles b.80 miles c.100 miles d.120 miles

OpenStudy (_.mufasa._):

C. 100 miles

OpenStudy (anonymous):

a scale drawing of a spaceship has a height of 5.5 inches. If the drawing has a scale of 1 inch=34 feet, what is the actual height of the spaceship? a.39.5 ft b.177 ft c.187 in. d.187 ft.

OpenStudy (_.mufasa._):

D. 187 feet

OpenStudy (anonymous):

a model sailboat is built using a scale of 1 cm= 12 ft. If the actual length of the boat is 84 ft, how long is the model boat? a.7 cm b.6 cm c.8 ft d.7 ft

OpenStudy (_.mufasa._):

A. 7 cm

OpenStudy (anonymous):

How long does it take for you to just learn the concept yourself? Look at why the answers above are correct, and actually learn the concept.